OpenStudy (anonymous):

defined function g(x) = \[\ln ( x ^{2} + 1)\] primitive calculates g whose graph passes through the origin of coordinates

terenzreignz (terenzreignz):

I'm sorry, what was the question? :3

OpenStudy (anonymous):

\[\int\limits\ln ( x ^{2} + 1)\]

OpenStudy (dumbcow):

ok start with integrating by parts u = ln(x^2 +1) dv = dx du = 2x/(x^2 +1) v = x

OpenStudy (anonymous):